Solve by elimination or substitution method
5x-2y = 23
2x+3y =13

Answer by stanbon(75887)

Solve by elimination or substitution method
5x-2y = 23
2x+3y =13
---
Multiply thru 1st by 3
Multiply thru 2nd by 2
----
15x-6y = 3*23
4x +6y = 2*13
--------------------
Add and solve for "y":
19x = 95
x = 5
------------
Solve for "y":
5x-2y = 23
5*5 - 2y = 23
25 -2y = 23
-2y = -2
y = 1
==============
Cheers,
Stan H.
===========